You suspect Ford might be undervalued. Use Bloomberg to get Ford’s EBITDA for “2017 Y 12/31/2017” and GM’s enterprise value/EBITDA ratio to calculate an enterprise value for Ford. To get Ford’s EBITDA, use the Financial Analysis tool, go to I/S, then select Adjusted. Scroll to the bottom of the page. To get GM’s enterprise value and EBITDA (both adjusted), use the financial analysis tab and go to the Key Stats page. You will need the “2017 Y 12/31/2017” numbers.
Answer #1

Using Bloomberg:

Ford EBITDA for year ended 31/12/2017 = $ 13623 million

GM's EBITDA = $ 21870 million and GM's EV (Enterprise Value) = $ 52550

GM's EV/EBITDA = 52550 / 21870 = 2.402835

Ford's EV = Ford's EBITDA x GM's EV/EBITDA = 13623 x 2.402835 = $ 32733.82 million